09 April 2014 DD 1/2 Night

Mostly clear with some cirrus.

E-time first half of night. See Steve's notes - lots of issues.

Tested Michele de la Pena's patch to PCS to allow NS tracking at a specific latitude/longitude of a NS target.
  • Worked.
  • In NSIGUI, used target name g: 270.1, -25, 0 @ 301 and clicked Get Ephemeris, then Set. Then I sent the mods acquisition script, moonS1.acq (in /home/lbto/SciPrep/MODS/Moon).

Tested large relative DETXY offsets during this NS track preset (as sent from the MODS UI, but these can be written into the script).
  • Worked.
  • However, note that TELRA, TELDEC do not reflect these offsets (it seems that telra, teldec reflect radec offsets only, not detxy offsets). No header info on the translational offsets. POSANGLE does show the new position angle.
    • MODS images mods1r.20140409.0015-17.fits are at specified position (15), detxy offset: offsetxy 90 1000 900 rel (16) and detxy offset: offsetxy -90 -1000 -900 rel (17).

07:50 acqMODS J1139+6353_p8asec_HA2p5.acq PA=-55 and using 0.8" slit since seeing is so good and moon is still up (but nearly set). GS is faint (R=15.53). Changing the slit to 1" since seeing is not so good and changed GS to a brighter one (R15.26).

08:05 acqMODS J1139+6353_1asec_HA2p5.acq (ERROR: red GO exposure already in progress - this was after I'd cntl'cd the previous script run with acqMODS -a).

08:10 Take confirm image through 1.2" slit (seeing on guider is up to 1.4" now), but WFS FWHM is 0.9" still. Confirmatory image through 1.0" slit is red 23, through 1.2" slit, red 24.

08:15 execMODS J1139+6353_1asec.obs airmass 1.29 (but cntl'ced twice as setting up - seeing is on the edge).

08:19 execMODS J1139+6353_1asec.obs
  • SNR is very low even in these better conditions. Cirrus present, about 0.5 mag extinction dips visible on guide plot. Oddly, the R magnitude of the guide star as showed on LBTplot is 20.0, but the star is 15.26. The exposure time is 2-sec for the guider images. The total integrated flux shown by GCS is 1700. Why so low? I used ND1.5 earlier for the moon observation, but every indication from the MODS logs and gui show that the clear filter is now in place.

09:15 acqMODS HIP55974_1asec.acq V=8.1 A3V. Very much saturated on 0.5-sec acq image through r-sdss, and slit hardly visible. modsAlign to posn star at slit X-center and Y=551, followed by DETXY offset +4.8". No confirmation, just observe spectrum. 4-sec exposure time gives peak counts of 4000 in blue and 10000 in red.

09:30 There is really no time to set up and observe a flux calibrator at 5" and 1" slits for a measure of slit losses. It takes ~25 minutes to switch to LBCs. The WFS FWHM was hovering around 1". I slewed to HZ44, but, realizing it is in quite a different part of the sky (2 hrs in RA from the qso), I wondered whether it would have given much indication of the slit losses for the qso. The J1139 spectra are disappointingly faint (in retrospect, should not have tried this faint an object, g=20.3, with some moon). With the more cirrusy data from 20140124, we have a total now of 5400 sec on this quasar.

09:40 Steve reconfiguring for the LBCs. John creating the script to run the set of 3. Before sunrise, we'll try to obtain the flats for this program (g-SLOAN in B and r-SLOAN in R and for Zheng Cai's program (Uspec in B and i-SLOAN in R).

10:07 Collimate - Steve needed to send Z8 = +2000 nm manual correction.

10:15 Copoint

10:19 Collimate on first science field.

10:23 Launch series of OBs via John's script loop_observe.py in ~lbcobs/supportscripts/SymbioticSupport/

10:36 First series of 3 shows strong elongation on Blue --- stop to check focus/collimation now. Also edited the script to use the corrected B1 ob (with type filled in as "object" and not blank). Missing type was causing some weirdness in UI, may be implications in header values, though RA and DEC at least are correct and OB played as expected.

10:44 Still a sign of elongation in the blue. There was a strong gradient from Saturn on the first field on which we focussed/collimated. Might have been better to focus/collimate on a field with no background gradient. But the field is setting - we'll continue. Blue rotator stopped working. Error message in LBC log. Stopped script and now taking a test exposure at PA=20 (was at PA=0), but the blue rotator is not going to position.

10:57 Cycling power to other systems. This seems to have resolved the blue rotator problem. Playing script again. Red guide stars appear round and FWHM 0.9", but Blue guide stars remain elongated (and I expect the images will follow suit).

11:10 Script was only repeating the first two - John fixed this. Stopped script and restarted on 3rd OB with "-f 3" flag. Now we are getting all three.
  • There are numerous LBC B warnings about the Rotator and FW #1. The FW #1 error message continued after cycling power, but FW #1 seems to recover when a new OB is loaded and played (as noted in an IT 5065).

12:00 Sky background is increasing - to near saturation in the blue. Going to stop this and slew to the flat field position.

12:19 Starting flats at g/i (B/R). PA=180. wfsblank1 at HA=1, DEC=73.
  • PA=180 3-sec some stars visible still 13k --- 21k
  • PA=0 1-sec 14k-28k

12:29
  • Flats at S/r (B/R).
  • PA=0 1-sec... 30k/24k --- 51k/48k --- 61k
  • PA=180 0.7sec ... takes a while for the derotator... sky too bright now. Stopped OB.

12:37 Steve closing up.

12:51 Calibrations:
  • MODS
    • grpixflats.cal
    • grlamps.cal
  • LBCs 25-biases x 2

Summary: On the engineerng half, the first task was tried but an error was encountered and no more could be done. This freed some time to observe an SDSS quasar which the ASASSN survey had noticed had brightened by 3 magnitudes. About ~1 hr was spent on this (for detailed times, see OSA log). Then many problems were encountered with multiple TCS systems. When LSS was stopped, the problems went away. Tried to reproduce the error, but I think with no success. Tested NS tracking on a part of a NS target - worked fine. Tested large (~1500" and 90 deg) relative offsets from MODS - worked fine. Juan Carlos obtained some data to compare DIMM vs GCS seeing.

On the D-half: Switched to MODS and tried to re-observe a quasar that was observed through cirrus in January. Guide star of R=15.2 appeared very faint and the LBTplot showed its magnitude to be 20th. Exposure time for guide camera was 2-sec, which is what LBTplot assumes. Integrated flux on GCS guide was extremely low, ~1700. Why so low? I'd used the ND1 filter in the guider for the moon (we didn't guide, but just in case), but there was no indication that it was still in place. Guide star for next object was OK. Made it hard to estimate seeing intially, although from the acq images and WFS FWHM, seeing was <~1". Decided not to use any larger slit. But decision to observe the fainter of the 2 quasars, at g=20.3, was not a good one and we did not add much to the January data. The sky was just too bright. Sigh...

Second part of D-half: Switched to the LBCs. DX required 2000nm Z8 for dofpia to work on pupils. John wrote a wrapper to send the 3 OBs one after another --- some editing of this was needed, but finally OK. Blue images are quite elongated, though red images look OK. We decided to keep going to collect enough data at least on red. Had a Blue rotator error (IT 5132) for which we turned off/on systems to recover and the now-becoming-familiar FW #1 timeout error (IT 5065) which was resolved upon loading/playing a new OB.

All-in-all a "learning experience"...

-- %USERSIG{OlgaKuhn - 2014-04-09}%

Comments

 
Topic revision: r4 - 09 Apr 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback